OpenStudy (anonymous):

Find the exact value of tan(x-y) if sin x = 8/17 and cos y = 3/5 such that 0

OpenStudy (anonymous):

0 < x < pi/2 and 0 < y < pi/2 means they are first-quadrant angles. tan (x - y) = (tan x - tan y)/(1+ tany tan x)

OpenStudy (anonymous):

If sin x = 8/17, then tan x = 8/15, the 8-15-17 right triangle If cos y = 3/5, then tan y = 4/3, the 3-4-5 right triangle So use the values of tan x and tan y in the ormula, and you'll be done.

OpenStudy (anonymous):

Thank you again (: